OpenStudy (anonymous):

what is the slope of a line given by the equation y=4x? enter your answer as an integer or fraction in lowest terms

OpenStudy (anonymous):

@blurbendy

jimthompson5910 (jim_thompson5910):

Hint: y = 4x is the same as y = 4x + 0 compare this with the form y = mx+b

jimthompson5910 (jim_thompson5910):

with y = mx+b, m is the slope, b is the y-intercept